OpenStudy (ttop0816):

am i correct? Which of the following shows the correct order for -1 2/5, (-1)^2, -1.4 and (1/2)2? (-1)^2 < -1.4 < -1 < (1/2)^2 -1.4 < -1 2/5 < (1/2)^2 < (-1)^2 << -1 2/5 < -1.4 < (1/2)^2 < (-1)^2

OpenStudy (anonymous):

No option C is correct....

OpenStudy (anonymous):

becouse -12/5 ~ -2.4 < -1.4 and so on......

OpenStudy (anonymous):

Is it clear.... do you want full explanation.... @ttop0816

OpenStudy (ttop0816):

@chetan552 its clear! thank you (:

OpenStudy (anonymous):

You are always welcome....@ttop0816

OpenStudy (ttop0816):

@kc_kennylau do you mind checking if my answer choices are correct?? Which of the following cosine functions has a period of 3 π ? y = 3cosx y = cos3x << y = cosx y = cosx The period of y = sin x/4 is _____. 1/4 π <<< 1/2 π 2π 8π https://media.glynlyon.com/g_trg_2013/2/graph_secx_2.gif y = csc y = sec x/2 << y = csc2x y = sec2x

OpenStudy (ttop0816):

@tester97 mind helping out also? (:

OpenStudy (kc_kennylau):

1. nope :D

OpenStudy (kc_kennylau):

2. nope :D

OpenStudy (ttop0816):

);

OpenStudy (kc_kennylau):

3. nope :D

OpenStudy (ttop0816):

@kc_kennylau oh! then would the first question be y = cos2/3x??

OpenStudy (kc_kennylau):

1. yes :D

OpenStudy (ttop0816):

the second question would be 8pi??

OpenStudy (kc_kennylau):

2. yes :D

OpenStudy (kc_kennylau):

Look at the graphs of each trigonometric functions first: http://img.sparknotes.com/figures/A/ad79275cb59e569b790cb945a4ffc553/triggraphs.gif

OpenStudy (kc_kennylau):

Which does the given graph look like?

OpenStudy (ttop0816):

its looks like a sec graph?

OpenStudy (kc_kennylau):

omg you're correct i'm so sorry :/